# Gen-T: Table Reclamation in Data Lakes

### 21 Mar 2024 v1

@@ I Introduction
Suppose a user is reading a news article that reports the demographics of employees in Top US tech companies in 2021 (top blue table in Figure 1). The user has access to a 2021 Microsoft’s Diversity Report [1] which seems to contradict the numbers in the news article (bottom green table in Figure 1). Using table reclamation and her data lake (which may include wikitables, NYTimes Data and other public and private datasets), the user can ask if there is a set of tables that, when integrated, recreates the data in the news article (blue table). Table reclamation is able to reproduce this table using a number of tables including World_MS_Ethnicity and World_MS_Employees (which, after being joined and a selection on 2021 is applied, produce the first tuple). We call the tables used in reclamation, originating tables. Other tables can be unioned with the Microsoft tables to reclaim the other tuples. From this (the originating tables including their meta-data and data), a user can understand that while her table (green) is reporting US statistics numbers, the article is reporting international numbers.
@@ IV-B Preliminaries: Integration Operators
with the same schema, complements if they share at least one non-null column value a less restrictive condition than subsumption, and contains some non-null values where has nulls while contains some non-null values where has nulls. The tuples must agree on all values on which they are both non-null.
@@
Given two tables that contain no duplicate tuples, and no tuples that can be subsumed or complemented, for all SPJU queries, there exists an equivalent query consisting of only Outer Union and the four unary operators (selection, projection, complementation, and subsumption).222The proof is included in Appendix -A
@@ V-A1 Candidate Table Retrieval
With a set of tables returned as relevant to the Source Table, we need to verify the set similarities of their values with the Source Table. To do so, we retrieve candidate tables among the previously discovered tables using a set similarity algorithm. This could be done efficiently with a system like JOSIE [32] that computes exact set containment or MATE [36] that supports multi-attribute joins. In addition to finding candidate tables containing columns that have high set similarity with a Source Table, we also diversify the set of candidate tables such that each candidate table has minimal overlap with other candidates. This is especially important in public data lakes, which tend to have multiple versions of the same tables [62, 74] and a large percentage of duplicate column sets [32]. By diversifying candidates, each candidate may overlap with different values in , as illustrated in the following example.
@@
Suppose we have the Source Table from Figure 3. In addition to data lake tables A, B, C, D, we also have Table E, an exact duplicate of Table D. If we only rank these tables using set overlap with the Source Table, Tables D and (its duplicate) Table E become top candidates, since all their columns have high set overlap with those in the Source Table. However, Table E does not add any new information when integrated with Table D. Thus, diversifying the set of candidate tables decreases Table E’s score, pushing other tables such as Table A higher in the ranking.
@@ V-A2 Matrix Traversal
First, we need to align tuples in candidate tables to tuples in a Source Table, based on shared values with the key attribute from the Source Table. To do so, we need to ensure that each candidate table contains a key column of the Source Table. If it does not, we greedily find a best way to join it with candidates that include the source key. We use standard join cardinality estimation to find a path that covers the most source key values and is as close to functional as possible (this procedure is denoted as Expand()).333The Expand algorithm is included in Appendix A-A This way, all tables can align its tuples with the Source Table using key values.
@@ VI-A Experimental Setup
TP-TR Benchmarks: We use the 8 tables from the TPC-H benchmark [76], which contain business information including customers, products, suppliers, nations, etc. Using these tables, we create three versions of a benchmark suite titled TP-Table Reclamation (TP-TR). TP-TR Large has TPC-H tables with original table sizes. TP-TR Med has TPC-H tables that are each 1/100 of its original table’s rows, and TP-TR Small has TPC-H tables that are each 1/1000 of its original table’s rows. In each, we take each of the 8 tables and create 4 versions of the same table – creating 32 tables in total. For two versions, we randomly nullify different subsets of values, and for the other two versions, we randomly inject different non-null (erroneous) values in different subsets of values. For the majority of the experiments, the number of nulls (respectively, erroneous values) is 50% meaning that we randomly take 50% of each table’s values and replace them with nulls (respectively, replace them with different new strings). In an ablation study (Section VI-B last two paragraphs), we vary (independently) the number of nulls/erroneous values from 10 to 90%. Our goal in the table discovery phase is then to filter out originating tables with injected non-null noise, so that resulting reclaimed tables do not contain erroneous values. Thus, we seek to verify that our approach uses the nullified versions rather than the erroneous versions so that combining them can reproduce the Source Table.
@@ VI-A1 Baselines
Auto-Pipeline has a similar framework to our problem in discovering the integration (query or pipeline) that reclaims a Source Table, however Gen-T does not assume to have the perfect set of input tables from which we can synthesize the query that reproduces the Source Table. Auto-Pipeline has both query-search and deep reinforcement learning approaches, but since we propose an unsupervised approach, we use the query-search variation as our baseline. Auto-Pipeline’s code implementation is not openly available, so we adopted an open re-implementation of their search approach [62], which adapts the framework in Foofah [80], and, for a fair comparison, revised their set of table operators to only contain table operators that Gen-T considers ({}). We call this re-implemented, adapted baseline Auto-Pipeline*. Since Auto-Pipeline’s benchmarks contain small tables, and most of their operators are string-transformation operators, we do not consider their benchmarks for our experiments.
@@ VI-A2 Metrics
Conditional KL-divergence: We aim to capture how many erroneous values there are in aligned tuples from a reclaimed table with respect to tuples in a Source Table. We adopt the traditional definition of conditional KL-divergence [81, 82], and add a penalization for erroneous values, such that the score is higher (diverges more) for reclaimed tables containing erroneous values as opposed to nulls in their aligned tuples with the Source Table (see Appendix A-C for more details). We report average scores over all Source Tables.
@@ VI-B Effectiveness
Across all benchmarks, Gen-T outperforms the baselines for all metrics, while perfectly reclaiming 15-17 Source Tables across all benchmarks. The baselines ALITE-PS and Auto-Pipeline* only perfectly reclaim 3 Source Tables and 1 Source Table across the benchmarks on which they do not time out, respectively, and ALITE does not perfectly reclaim any. In fairness, ALITE is an integration method and does not take the Source Table into account (it is not ”target-driven” like Auto-Pipeline*). In terms of similarity (Table II), Gen-T outperforms the top performing existing baseline method (ALITE-PS) by 11-25% in Recall and by 48-56% in Precision across all TP-TR benchmarks. For the divergence measures (bottom table of Table II), we see that Gen-T produces tables that contain fewer nullified values in its aligned tuples with respect to the Source Table (Inst-Div.), as well as fewer erroneous values in its aligned tuples, which is reflected in the lower D scores than the baselines.
@@ VI-D Generalizability
We then run experiments on a data lake consisting of both T2D Gold and WDC Sample tables. This way, we evaluate how well the methods perform when a set of candidate tables returned from Set Similarity may contain irrelevant or misleading tables from the WDC Sample benchmark. Table IV presents the similarity and divergence scores for all methods on 33 of the common sources from T2D Gold for which all methods have non-empty, reasonably sized output tables. We can see that Gen-T outperforms the baselines for all measures, even having perfect precision of 1.0. In contrast, the baseline methods that are given the candidate tables from Set Similarity integrate all candidate tables and produce tables that contain a lot of additional tuples.
@@ VII Conclusion and Open Problems
Table Reclamation is essential in verifying if a data lake supports the tuples (facts) in a Source table. Our results show that, despite the large search space, Gen-T can solve the reclamation problem efficiently for source tables with keys. In future work, we will relax the key assumption with regard to source tables, and use a fast, approximate instance comparison algorithm to compare instances from a source table and data lake tables [83]. When a table can only partially be reclaimed, we plan to investigate whether the originating tables can be embedded in a new data lake and used to possibly generate a better reclamation. Alternatively, we plan to investigate if reclamation can be combined with data cleaning (for example, value imputation over missing values or entity resolution) to produce a better reclamation. In addition, we plan to consider the case in which values from a source table do not syntactically align with values from a data lake, in which case we can explore the semantic similarity of instances. Table reclamation can also be used to verify the tabular results of generative AI or large language models. Verifying the output of generative AI using a data management lens is an emerging and important area [84]. For example, users who generate summary tables and charts (e.g. Microsoft Copilot [85]) or presentation slides (e.g. SlidesAI [86]) from input data would find it useful to verify model outputs and examine what data was used to generate them. Our approach will advance this area by allowing the automatic verification of tables that are created through complex integrations of other tables, something not yet considered in the literature.

### 22 Mar 2024 v2

@@ I Introduction
Suppose a user is reading a news article that reports the demographics of employees in Top US tech companies in 2021 (top blue table in Figure 1). The user has access to 2021 Microsoft’s Diversity Report [1] which seems to contradict the numbers in the news article (bottom green table in Figure 1). Using table reclamation and her data lake (which may include wikitables, NYTimes Data and other public and private datasets), the user can ask if there is a set of tables that, when integrated, recreates the data in the news article (blue table). Table reclamation is able to reproduce this table using a number of tables including World_MS_Ethnicity and World_MS_Employees (which, after being joined and a selection on 2021 is applied, produce the first tuple). We call the tables used in reclamation, originating tables. Other tables can be unioned with the Microsoft tables to reclaim the other tuples. From this (the originating tables including their meta-data and data), a user can understand that while her table (green) is reporting US statistics numbers, the article is reporting international numbers.
@@ IV-B Preliminaries: Integration Operators
with the same schema, complements if they share at least one non-null column value, and contains some non-null values where has nulls while contains some non-null values where has nulls. The tuples must agree on all values on which they are both non-null.
@@
Given two tables that contain no duplicate tuples, and no tuples that can be subsumed or complemented, for all SPJU queries, there exists an equivalent query consisting of only Outer Union and the four unary operators (selection, projection, complementation, and subsumption).222The proof is included in our technical report [73].
@@ V-A1 Candidate Table Retrieval
With a set of top- tables returned as relevant to the Source Table, we need to verify the set similarities of their values with the Source Table. To do so, we retrieve candidate tables among the previously discovered tables using a set similarity algorithm. This could be done efficiently with a system like JOSIE [32] that computes exact set containment or MATE [36] that supports multi-attribute joins. In addition to finding candidate tables containing columns that have high set similarity with a Source Table, we also diversify the set of candidate tables such that each candidate table has minimal overlap with other candidates. This is especially important in public data lakes, which tend to have multiple versions of the same tables [62, 75] and a large percentage of duplicate column sets [32]. By diversifying candidates, each candidate may overlap with different values in , as illustrated in the following example.
@@
Suppose we have the Source Table from Figure 3. In addition to data lake tables A, B, C, D, we also have Table E, an exact duplicate of Table D. If we only rank these tables using set overlap with the Source Table, Tables D and (its duplicate) Table E become top candidates, since all their columns have high set overlap with those in the Source Table. However, Table E does not add any new information when integrated with Table D. Thus, diversifying the set of candidate tables decreases Table E’s score, pushing other tables such as Table A higher in the top- ranking.
@@ V-A2 Matrix Traversal
First, we need to align tuples in candidate tables to tuples in a Source Table, based on shared values with the key attribute from the Source Table. To do so, we need to ensure that each candidate table contains a key column of the Source Table. If it does not, we greedily find a best way to join it with candidates that include the source key. We use standard join cardinality estimation to find a path that covers the most source key values and is as close to functional as possible (this procedure is denoted as Expand()).333The Expand algorithm is included in the technical report [73]. This way, all tables can align its tuples with the Source Table using key values.
@@ VI-A Experimental Setup
TP-TR Benchmarks: We use the 8 tables from the TPC-H benchmark [77], which contain business information including customers, products, suppliers, nations, etc. Using these tables, we create three versions of a benchmark suite titled TP-Table Reclamation (TP-TR). TP-TR Large has TPC-H tables with original table sizes. TP-TR Med has TPC-H tables that are each 1/100 of its original table’s rows, and TP-TR Small has TPC-H tables that are each 1/1000 of its original table’s rows. In each, we take each of the 8 tables and create 4 versions of the same table – creating 32 tables in total. For two versions, we randomly nullify different subsets of values, and for the other two versions, we randomly inject different non-null (erroneous) values in different subsets of values. For the majority of the experiments, the number of nulls (respectively, erroneous values) is 50% meaning that we randomly take 50% of each table’s values and replace them with nulls (respectively, replace them with different new strings). In an ablation study (Section VI-B last two paragraphs), we vary (independently) the number of nulls/erroneous values from 10% to 90%. Our goal in the table discovery phase is then to filter out originating tables with injected non-null noise, so that resulting reclaimed tables do not contain erroneous values. Thus, we seek to verify that our approach uses the nullified versions rather than the erroneous versions so that combining them can reproduce the Source Table.
@@ VI-A1 Baselines
Auto-Pipeline has a similar framework to our problem in discovering the integration (query or pipeline) that reclaims a Source Table. However, Gen-T does not assume to have the perfect set of input tables from which we can synthesize the query that reproduces the Source Table. Auto-Pipeline has both query-search and deep reinforcement learning approaches, but since we propose an unsupervised approach, we use the query-search variation as our baseline. Auto-Pipeline’s code implementation is not openly available, so we adopted an open re-implementation of their search approach [62], which adapts the framework in Foofah [81], and, for a fair comparison, revised their set of table operators to only contain table operators that Gen-T considers ({}). We call this re-implemented, adapted baseline Auto-Pipeline*. Since Auto-Pipeline’s benchmarks contain small tables, and most of their operators are string-transformation operators, we do not consider their benchmarks for our experiments.
@@ VI-A2 Metrics
Conditional KL-divergence: We aim to capture how many erroneous values there are in aligned tuples from a reclaimed table with respect to tuples in a Source Table. We adopt the traditional definition of conditional KL-divergence [82, 83], and add a penalization for erroneous values, such that the score is higher (diverges more) for reclaimed tables containing erroneous values as opposed to nulls in their aligned tuples with the Source Table (see the technical report [73] for more details). We report average scores over all Source Tables.
@@ VI-B Effectiveness
Across all benchmarks, Gen-T outperforms the baselines for all metrics, while perfectly reclaiming 15-17 Source Tables across all benchmarks. Baselines ALITE-PS and Auto-Pipeline* only perfectly reclaim 3 Source Tables and 1 Source Table, respectively, across the benchmarks on which they do not time out, and ALITE and Ver does not perfectly reclaim any. In fairness, ALITE is an integration method that does not consider the Source Table (it is not ”target-driven” like Auto-Pipeline*). Also, Ver is a QBE method whose goal is to produce source tuples along with many additional tuples. In terms of similarity (Table II), Gen-T outperforms the top performing existing baseline method (ALITE-PS) by 11-25% in Recall and by 48-56% in Precision across all TP-TR benchmarks. For the divergence measures, we see that Gen-T produces tables that contain fewer nullified values in its aligned tuples with respect to the Source Table (Inst-Div.), as well as fewer erroneous values in its aligned tuples, which is reflected in the lower D scores than the baselines.
@@ VI-D Generalizability
We then run experiments on a data lake consisting of both T2D Gold and WDC Sample tables. This way, we evaluate how well the methods perform when a set of candidate tables returned from Set Similarity may contain irrelevant or misleading tables from the WDC Sample benchmark. Table IV presents the similarity and divergence scores for all methods on 33 of the common sources from T2D Gold for which all methods have non-empty, reasonably sized output tables. We can see that Gen-T outperforms the baselines for all measures, even having a precision of 1.0. In contrast, the baseline methods that are given the candidate tables from Set Similarity integrate all candidate tables and produce tables that contain many additional tuples.
@@ VII Conclusion and Open Problems
Table Reclamation is essential in verifying if a data lake supports the tuples (facts) in a Source table. Our results show that, despite the large search space, Gen-T can solve the reclamation problem efficiently for source tables with keys. In future work, we will relax the key assumption with regard to source tables, and use a fast, approximate instance comparison algorithm to compare instances from a source table and data lake tables [84]. When a table can only be partially reclaimed, we plan to investigate whether the originating tables can be embedded in a new data lake and used to possibly generate a better reclamation. Alternatively, we plan to investigate if reclamation can be combined with data cleaning (for example, value imputation over missing values or entity resolution) to produce a better reclamation. In addition, we plan to consider the case in which values from a source table do not syntactically align with values from a data lake, in which case we can explore the semantic similarity of instances. Table reclamation can also be used to verify the tabular results of generative AI or large language models. Verifying the output of generative AI using a data management lens is an emerging and important area [85, 86]. For example, users who generate summary tables and charts (e.g. Microsoft Copilot [87]) or presentation slides (e.g. SlidesAI [88]) from input data would find it useful to verify model outputs and examine what data was used to generate them. Our approach will advance this area by allowing the automatic verification of tables that are created through complex integrations of other tables, something not yet considered in the literature.